OpenStudy (pandaecat):

A rectangle has a length of the cube root of 81 inches and a width of 3 to the 2 over 3 power inches. Find the area of the rectangle. A 3 to the 2 over 3 power inches squared B 3 to the 8 over 3 power inches squared C 9 inches squared D9 to the 2 over 3 power inches squared I think it's C but I'm not really sure. Can someone explain to me?

OpenStudy (egbeach):

the formula for area is length x width

OpenStudy (pandaecat):

@egbeach I know that... but I just don't get how u simply the thing

OpenStudy (egbeach):

\[\sqrt[3]{81} \times 3^{\frac{ 2 }{ 3 }}\]

OpenStudy (egbeach):

thats the equation right?

OpenStudy (pandaecat):

Yes

OpenStudy (egbeach):

well whats \[\sqrt[3]{81}\]

OpenStudy (pandaecat):

81 and one third

OpenStudy (egbeach):

\[81^{\frac{ 1 }{ 3}}\times 3^{\frac{ 2 }{ 3}}\]

OpenStudy (egbeach):

are you able to solve that?

OpenStudy (pandaecat):

Um er a moment

OpenStudy (egbeach):

it should work if you plug it into a calculator correctly

OpenStudy (pandaecat):

54???

OpenStudy (egbeach):

ummm no... the answer is actually 9 so you were right

OpenStudy (pandaecat):

Oh then... I'll check again... Just confused

OpenStudy (egbeach):

ok